# Waver: Wave Your Way to Lifelike Video Generation

### v1

@@ 1 Introduction
To overcome these problems and challenges, we propose Waver, a model based on Rectified Flow Transformer (Esser et al., 2024b) and meticulously designed to achieve industry-grade performance. Waver consists of two modules: Task-Unified DiT and Cascade Refiner. We first employ Task-Unified DiT to generate videos with a resolution of 720p and durations ranging from 5 to 10 seconds, and subsequently apply Cascade Refiner to upscale them to 1080p resolution.
@@ 2 Model Architecture
We use Wan2.1-VAE (Wan et al., 2025) for efficient video latent compression and a powerful dual-encoder system, combining flan-t5-xxl (Raffel et al., 2020) and Qwen2.5-32B-Instruct (QwenTeam, 2024), for superior text understanding. Benefiting from the strong performance of Qwen, this dual-encoder setup demonstrably improves prompt following in the text-to-image (T2I) task over the single-encoder (flan-t5-xxl) baseline (Fig. 4). The architecture of Waver primarily consists of two components: the Task-Unified DiT and the Cascade Refiner. The Task-Unified DiT is built upon rectified flow Transformers (Esser et al., 2024b), serves as the core generation model. It fuses video and text modalities through a ”Dual Stream + Single Stream” architecture. The Cascade Refiner is also based on rectified flow Transformers, functions as a super-resolution module. It takes the 480p or 720p videos generated by the Task-Unified DiT as input and upscales them to a final resolution of 1080p. This design enables efficient, high-fidelity video super-resolution with fewer inference steps.
@@ 3.3 Caption Model
Model Training. Based on the Qwen2.5-VL (Bai et al., 2025), we conduct joint training with both captioning and sub-action description tasks. For the captioning task, the model takes the Tariser2 (Yuan et al., 2025b) caption as input and generates detailed descriptive captions as output. The visual encoder is kept unfrozen during this stage. The training conducted during the DPO stage is aimed at mitigating hallucination. On the DPO (Rafailov et al., 2024) side, DPOP (Pal et al., 2024) loss is employed to maintain the stability of long-form outputs. In this stage, the visual encoder is frozen. The captioning model is trained on English-language data.
@@ 4.1 Multi-stage Training
Training hyperparameters. Throughout the entire training process, we ensure that the data from each stage is seen for at least one epoch, and the overall learning rate shows a decreasing trend. Following SD3 (Esser et al., 2024a), the sigma shift in flow matching increases as the resolution increases.
@@ 4.3 Motion Optimization
Noise scheduling. We follow the flow matching schedule proposed in SD3 (Esser et al., 2024a). It highlights that timestep sampling is crucial for flow matching training. Specifically, timesteps near the middle of the [0, 1] interval are more challenging for the model to learn. Therefore, probability density functions that assign higher probability to the middle and lower probability to the ends, such as the logit-normal and mode functions, are used for timestep sampling, as is shown in Eq. 1 and Eq. 2, respectively.
@@ 7.1 DiT Model Sparsity
These observations suggest that a sparse attention strategy adaptive to the intrinsic structure of attention patterns would outperform fixed-window approaches. Based on this insight, we explore two potential directions: 1) Spatial-Temporal Sliding Window Attention. This method offers better alignment with spatio-temporally local attention patterns compared to conventional window attention. Although it has been explored in prior work (Xi et al., 2025; Zhang et al., 2025b), it remains limited in flexibility and is ineffective at capturing non-local dependencies such as fixed-column patterns. 2) NSA-Based Adaptive Sparse Attention. Native Sparse Attention (NSA) (Yuan et al., 2025a) provides greater flexibility in designing attention patterns. However, existing NSA methods are primarily designed for one-dimensional text data. Effectively adapting them to video—a 3D spatio-temporal medium—requires careful redesign to incorporate its structural properties.
Building on these findings, our future work will focus on developing advanced sparse attention mechanisms that not only enhance computational efficiency, but also enable high-performance real-time video generation, thereby paving the way for broader practical adoption.
(a) An overview of our VAE and ShuffleDown/Up blocks.
(b) Comparison of the text fidelity between Wan 2.1 VAE and our VAE on I2V generation task.
Figure 30: Overview of our proposed video VAE and its effectiveness in improving text fidelity.
7.2 VAE Impact on Video Generation
Video VAE not only encodes videos into latent tokens to accelerate diffusion training, but also has a significant impact on video generation quality. In this section, we introduce our VAE training recipes for addressing visual quality issues in video generation, including text fidelity, grainy textures, and background distortion.
Improving Text Fidelity with ShuffleDown/Up Blocks
We build a causal video VAE that compresses the temporal dimension by 4 times and each spatial dimension by 8 times. Different from previous works (Yang et al., 2025; Wan et al., 2025), we introduce pixel-shuffle and pixel-unshuffle operations (Shi et al., 2016) to develop the 3D downsample blocks and 3D upsample blocks. During downsampling, pixel unshuffle can directly map spatial information to the channel dimension and reduce information loss. During upsampling, pixel shuffle can better restore the spatial structure. An overview of our VAE is shown in Fig.30(a). The yellow block is the ShuffleDown block for downsampling, and the green block is the ShuffleUp block for upsampling. $T$, $H$, $W$, and $C$ indicates the input tensor shape. $C_{in}$ denotes the number of input channels of each block, and $C_{out}$ denotes the number of output channels of each block.
The proposed ShuffleDown and ShuffleUp blocks significantly enhance the video reconstruction capability of our VAE, thus improving the text fidelity of the I2V task. As shown in Fig. 30(b), the I2V model with our VAE preserves the text details in the conditional image, whereas the Wan 2.1 VAE loses these details and generates videos with distorted text.
Mitigating Grainy Textures and Background Distortion with KL Loss
@@ 7.2 VAE Impact on Video Generation
The KL divergence loss in our VAE is computed in closed form for the Gaussian case as follows:
@@
where $D$, $T$, $H$, and $W$ denote the latent dimension, temporal length, height, and width of the latent tokens, respectively. And $\mu_{dthw}$ and $\sigma_{dthw}$ are the mean and standard deviation at each position. Unlike (Esser et al., 2021), which computes the KL loss as a sum, we take the mean over all positions to ensure compatibility with multi-resolution training. The KL loss serves as a regularization objective, encouraging the learned latent distribution to approximate a standard normal distribution. The weight of the KL loss, denoted as $\lambda_{4}$ in Eq. 4, controls the strength of the regularization.
Our findings indicate that the KL loss in VAE training has a greater impact on the video generation task than on the VAE reconstruction task. The KL loss affects the generated video quality in two ways: (1) if the KL loss weight is too small, the decoder may generate grainy textures; (2) if the KL loss weight is too large, the background can become easily distorted. We investigate this problem within the I2V (Image-to-Video) task to enable precise control over the generated scenes We explore this problem in the context of the I2V (Image-to-Video) task, as I2V allows us to consistently control the generated scenes and ensure a fair comparison. Specifically, we finetune the Waver I2V model using different VAEs, and use the same conditional image and prompt as inputs across all I2V models. We then compare the videos generated by our VAE with those produced by the original Wan 2.1 VAE (Wan et al., 2025) and the CogvideoX VAE (Yang et al., 2025), as illustrated in Fig. 31.
The results indicate that when the KL loss weight is set too low, videos generated by the Waver I2V model with Wan VAE display obvious grainy textures. Note that the tokens predicted by the diffusion model are generally less accurate than those generated by the VAE encoder. Consequently, prediction errors may lead to mode collapse if the VAE decoder lacks robustness. A larger KL loss weight encourages VAE to predict higher values of $\sigma_{dthw}$, resulting in a noisier sampling process. The reparameterization-based sampling in the VAE acts as an augmentation mechanism for the decoder, thereby enhancing its robustness in video generation. In contrast, if the KL loss weight is too high, the Waver I2V model with CogvideoX VAE can cause the background to become obviously distorted in some cases. In the second row of Fig. 31, the ground appears swollen in the videos generated by CogvideoX VAE, whereas it is flat in the corresponding condition image.
Figure 31: Comparison of generated video quality with different VAEs on the I2V task.
Next, we examine how the KL loss weight influences the VAE latent distribution and, in turn, affects the learning of the diffusion model. We visualize the latent distribution of CogvideoX VAE and its KL=0 variant, as shown in Fig. 32. We find that even though the two variants of VAEs have similar reconstruction results, their latent distributions are largely different. When the KL loss weight is set to 0, the VAE is able to learn a smooth and well-structured latent space. In contrast, as the KL loss weight increases, the KL term begins to dominate the training objective. The learned latent representation begins to collapse to the prior distribution $\mathcal{N}(0,I)$. As a result, the latent representations become noisier, thereby increasing the learning difficulty for the diffusion model. Based on these observations, we finally set the KL loss weight to 3e43e-4 to train our VAE.
@@
Decrease LPIPS Loss Weight to Avoid Grid-like Artifacts
@@ 7.3 Caption Impact on Video Generation
The use of detailed caption models facilitates more precise alignment between video content and its corresponding captions, thereby enabling more effective training of Text-to-Video (T2V) models. This improved alignment during the pretraining phase contributes to a reduction in motion distortion. Furthermore, because detailed captions describe actions with greater specificity and adhere strictly to temporal order, T2V models exhibit enhanced instruction-following capabilities with respect to action sequences. In addition to actions, detailed captions provide richer descriptions of visual elements, such as the appearance and quantity of subjects, which further improves the model’s ability to generate visuals that are more consistent with real-world scenes. Fig. 34 illustrates the advantages of our T2V model trained with more detailed captions compared to that trained with Tarsier2 (Yuan et al., 2025b) captions.
Due to the lack of explicit indication regarding whether spatial relationships are described from the viewer’s perspective or from the perspective of the subjects within the video in publicly available Multimodal Large Language Model (MLLM) pretraining datasets, MLLMs often make errors when generating captions that reference spatial relations such as ”left” or ”right.” Specifically, the models may either incorrectly specify the spatial relation, omit the relevant perspective, or entirely fail to describe the spatial relationship. This ambiguity negatively impacts the alignment between the video content and the generated caption.
To address this issue in subsequent optimization of the captioning model, we place particular emphasis on the accurate representation of spatial relationships. We construct a comprehensive set of question-answer pairs focused on spatial relationships, with each question explicitly requiring temporal/sequential localization, subject/object localization, and action localization. This design ensures that each question refers to a unique event or moment within the video.
Furthermore, we increase the complexity of sub-action description task by incorporating queries spanning multiple temporal intervals and expand the volume of caption training data accordingly. The three tasks are jointed trained. In the evaluation conducted on the sports caption benchmark dataset, the proposed model demonstrated notable improvements over the Tarsier2 model across three key metrics. Specifically, the new model achieves more accurate and detailed descriptions, exhibits less action description hallucinations, and provides more precise spatial relation descriptions. The win rates (GSB) of these three indicators are 53.4%, 20.5%, and 39.2%.
@@ 9 Contributions and Acknowledgments
Yifu Zhang, Hao Yang, Yuqi Zhang, Yifei Hu, Fengda Zhu, Chuang Lin, Xiaofeng Mei
Yi Jiang, Zehuan Yuan, Bingyue Peng
@@ Acknowledgments:
Fangzhou Ai, Jinlai Liu, Xugang Ye, Xiaoran Xu, Bobo Zeng, Aaron Shen, Mao Hu, Xiaoxiao Qin, Tingxuan Li, Wanxing Wang, Puke Zhang, Yufei Wu, Ruoyu Guo, Ge Bai, Dongyang Wang, Tiger Li, Shoufa Chen, Chongjian Ge, Shilong Zhang, Shukai Wang, Dingyuan Xu, Chetan Velivela, Kairong Sun, Kaihua Jiang, Junru Zheng

### v2

@@ 1 Introduction
To overcome these problems and challenges, we propose Waver, a model based on Rectified Flow Transformer (Esser et al., 2024) and meticulously designed to achieve industry-grade performance. Waver consists of two modules: Task-Unified DiT and Cascade Refiner. We first employ Task-Unified DiT to generate videos with a resolution of 720p and durations ranging from 5 to 10 seconds, and subsequently apply Cascade Refiner to upscale them to 1080p resolution.
@@ 2 Model Architecture
We use Wan2.1-VAE (Wan et al., 2025) for efficient video latent compression and a powerful dual-encoder system, combining flan-t5-xxl (Raffel et al., 2020) and Qwen2.5-32B-Instruct (QwenTeam, 2024), for superior text understanding. Benefiting from the strong performance of Qwen, this dual-encoder setup demonstrably improves prompt following in the text-to-image (T2I) task over the single-encoder (flan-t5-xxl) baseline (Fig. 4). The architecture of Waver primarily consists of two components: the Task-Unified DiT and the Cascade Refiner. The Task-Unified DiT is built upon rectified flow Transformers (Esser et al., 2024), serves as the core generation model. It fuses video and text modalities through a ”Dual Stream + Single Stream” architecture. The Cascade Refiner is also based on rectified flow Transformers, functions as a super-resolution module. It takes the 480p or 720p videos generated by the Task-Unified DiT as input and upscales them to a final resolution of 1080p. This design enables efficient, high-fidelity video super-resolution with fewer inference steps.
@@ 3.3 Caption Model
Model Training. Based on the Qwen2.5-VL (Bai et al., 2025), we conduct joint training with both caption and sub-action description tasks. For the caption task, the model takes the Tariser2 (Yuan et al., 2025b) caption as input and generates detailed descriptive caption as output. The visual encoder is kept unfrozen during this stage. The training conducted during the DPO stage is aimed at mitigating hallucination. On the DPO (Rafailov et al., 2024) side, DPOP (Pal et al., 2024) loss is employed to maintain the stability of long-form outputs. In this stage, the visual encoder is frozen. The caption model is trained on English-language data.
@@ 4.1 Multi-stage Training
Training hyperparameters. Throughout the entire training process, we ensure that the data from each stage is seen for at least one epoch, and the overall learning rate shows a decreasing trend. Following SD3 (Esser et al., 2024), the sigma shift in flow matching increases as the resolution increases.
@@ 4.3 Motion Optimization
Noise scheduling. We follow the flow matching schedule proposed in SD3 (Esser et al., 2024). It highlights that timestep sampling is crucial for flow matching training. Specifically, timesteps near the middle of the [0, 1] interval are more challenging for the model to learn. Therefore, probability density functions that assign higher probability to the middle and lower probability to the ends, such as the logit-normal and mode functions, are used for timestep sampling, as is shown in Eq. 1 and Eq. 2, respectively.
@@ 7.1 DiT Model Sparsity
These observations suggest that a sparse attention strategy adaptive to the intrinsic structure of attention patterns would outperform fixed-window approaches. In particular, two directions appear promising: 1) Spatial-Temporal Sliding Window Attention. This method offers better alignment with spatio-temporally local attention patterns compared to conventional window attention. Although it has been explored in prior work (Xi et al., 2025; Zhang et al., 2025b), it remains limited in flexibility and is ineffective at capturing non-local dependencies such as fixed-column patterns. 2) NSA-Based Adaptive Sparse Attention. Native Sparse Attention (NSA) (Yuan et al., 2025a) provides greater flexibility in designing attention patterns. However, existing NSA methods are primarily designed for one-dimensional text data. Effectively adapting them to video—a 3D spatio-temporal medium—requires careful redesign to incorporate its structural properties.
Overall, these findings underscore the importance of developing sparse attention mechanisms that are not only computationally efficient but also inherently adaptive to the multi-scale and dynamic nature of video. Such approaches are essential for enabling high-performance, real-time video generation and facilitating wider practical application.
7.2 VAE Empirical Findings on Video Generation
Given the critical role of Variational Autoencoders (VAEs) in video generation and their substantial influence on the training of diffusion transformers, our model, while leveraging the Wan 2.1 VAE, still necessitates a dedicated discussion on VAE optimization—given that VAE performance directly underpins video generation quality.
In this section, we elaborate on our VAE training protocols and empirical insights tailored to mitigating key visual quality challenges in video generation, such as perceptible grainy textures, unintended background distortions and grid-like artifacts.
KL Loss is Sensitive to Grainy Textures and Background Distortion
Figure 30: Comparison of generated video quality with different VAEs on the I2V task, where the VAEs are trained with different KL loss weights.
@@ 7.2 VAE Empirical Findings on Video Generation
The KL divergence loss in our exp is computed in closed form for the Gaussian case as follows:
@@
where $D$, $T$, $H$, and $W$ denote the latent dimension, temporal length, height, and width of the latent tokens, respectively. And $\mu_{dthw}$ and $\sigma_{dthw}$ are the mean and standard deviation at each position. The KL loss serves as a regularization objective, encouraging the learned latent distribution to approximate a standard normal distribution. The weight of the KL loss, denoted as $\lambda_{4}$ in Eq. 4, controls the strength of the regularization.
@@
We investigate this issue within the I2V task to ensure fair comparison through consistent scene control. By finetuning the Waver I2V model with different VAEs using identical conditional inputs, we compare videos generated by our VAE against those from Wan 2.1 VAE (Wan et al., 2025) and CogvideoX VAE (Yang et al., 2025) (Fig. 30). Our results indicate that a low KL loss weight (Wan VAE) causes grainy textures due to insufficient decoder robustness, which can lead to mode collapse. Conversely, an excessively high KL weight (CogvideoX VAE) distorts backgrounds. While a larger KL weight enhances decoder robustness through reparameterization-based sampling, finding an optimal balance is crucial.
We visualize the latent distributions of CogvideoX VAE and its KL=0 variant in Fig. 31. Although both VAEs achieve similar reconstruction, their latent spaces differ significantly. With KL loss weight at 0, the VAE learns a smooth, well-structured latent space. As the KL weight increases, the latent representation collapses toward the prior $\mathcal{N}(0,I)$, becoming noisier and making diffusion model training more challenging.
Without spatial relation QA pairs In an outdoor courtyard featuring dark tiled flooring, white ceramic planters shaped like toilets, and a backdrop of a wooden-and-metal fence adorned with hanging circular decorations, an older man (around 40 years old) stands near a house entrance. He wears a dark jacket, brown beanie, and blue jeans, holding something small in his hands as he looks toward the scene unfolding nearby. The camera slowly pans right, shifting focus from the man to a woman around 50 years old standing on a concrete path adjacent to the courtyard. She is dressed in a fluffy pink coat, black beanie, and red scarf, with one hand in her pocket and the other gesturing expressively—first waving, then pointing, and crossing arms—as she speaks. Two dogs accompany her: a brown one with a collar stands calmly beside her, while a white dog moves around her feet. The courtyard’s greenery, including yellow-leafed plants in the foreground and potted shrubs along the fence, frames the scene, with distant buildings and trees visible beyond. The woman continues her animated gestures.
With spatial relation QA pairs In an outdoor courtyard featuring dark tiled flooring, white ceramic planters shaped like toilets, and a wooden-and-metal fence adorned with hanging circular decorations, a man around 40 years old—dressed in a brown beanie, dark jacket, and blue jeans—stands near a house with a covered pergola and red-tiled roof, holding an object in his hands. The camera slowly pans right, shifting focus from the man to reveal a woman around 50 years old on the right side of the frame. She wears a pink fuzzy coat, black beanie, and red scarf, with her left hand in her coat pocket and her right hand making expressive gestures: first raising it with fingers spread, then pointing toward the camera, and finally crossing her arms. Beside her, two dogs—a brown one with a collar and a white one—stand on the concrete path, occasionally moving slightly. The background includes more potted plants, trees, and distant buildings with green fencing. As the camera continues panning, the man remains in the left background, still holding his object, while the woman stays centered, continuing her gestures and smiling. The dogs remain nearby, with the brown dog eventually turning its head toward the camera. The scene maintains a casual, warm atmosphere throughout, with the woman’s movements and the dogs’ presence adding dynamic elements to the static courtyard setting.
Table 4: Caption model trained without/with spatial relation QA pairs.
Decrease LPIPS Loss Weight Can Avoid Grid-like Artifacts
@@
Exploring Possible Future Directions for Video VAEs
(1) Higher Compression Ratios. Video VAEs with higher compression ratios are vital for efficient DiT training, as they reduce token length and enable longer video generation. The primary challenge is to advance this compression without degrading reconstruction and generation quality. (2) Fusion with Multimodality. Future Video VAEs should evolve from visual compressors into multimodal encoders. A key direction is to develop architectures that jointly embed video, audio, and text into a unified latent space, enabling more semantically rich and controllable multimodal generation.
@@ 7.3 Caption Impact on Video Generation
Detailed Caption is Indeed Important for Precise Instruction Following
The use of detailed caption models facilitates a more precise alignment between video content and its corresponding captions, enabling more effective training of Text-to-Video (T2V) models, which contributes to a reduction in motion distortion in the pretraining phase and improves the model’s ability to generate visuals that are more consistent with real world scenes. Furthermore, since detailed captions describe actions with greater specificity and adhere strictly to temporal order, T2V models exhibit enhanced instruction-following capabilities with respect to action sequences. Fig. 33 illustrates the advantages of our T2V model trained with more detailed captions compared to that trained with Tarsier2 (Yuan et al., 2025b) captions.
Enhancing Spatial Relation Representation is Helpful
Due to the lack of explicit indication regarding whether spatial relations are described from the viewer’s perspective or from the perspective of the subjects within the video in publicly available Multimodal Large Language Model (MLLM) pretraining datasets, MLLMs often make errors when generating captions that reference spatial relations such as “left” or “right”. Specifically, models may incorrectly specify the spatial relation, omit the relevant perspective, or completely fail to describe it. This ambiguity negatively impacts the alignment between the video content and the generated caption.
To address this issue in subsequent optimization of the caption model, we place particular emphasis on the accurate representation of spatial relations. We construct a comprehensive set of question-answer (QA) pairs focused on spatial relations, with each question explicitly requiring temporal/sequential localization, subject/object localization, and action localization. This design ensures that each question refers to a unique event or moment within the video. The cases of caption model trained without/with spatial relation QA pairs are presented in Tab. 4.
@@ 9 Contributions and Acknowledgments
Yifu Zhang, Hao Yang, Yuqi Zhang, Yifei Hu, Fengda Zhu, Chuang Lin, Xiaofeng Mei, Yi Jiang, Bingyue Peng, Zehuan Yuan
@@ Acknowledgments:
Fangzhou Ai, Jinlai Liu, Xugang Ye, Xiaoran Xu, Bobo Zeng, Aaron Shen, Mao Hu, Xiaoxiao Qin, Tingxuan Li, Wanxing Wang, Puke Zhang, Yufei Wu, Ruoyu Guo, Ge Bai, Xin Chen, Dongyang Wang, Tiger Li, Shoufa Chen, Chongjian Ge, Shilong Zhang, Shukai Wang, Dingyuan Xu, Chetan Velivela, Kairong Sun, Kaihua Jiang, Junru Zheng